OpenStudy (cookiimonster627):

solve each conditional equstion for x where 0 greater than or equal to x greater than 2pi. 3cos^2 x - sin^2 x =0

OpenStudy (dinamix):

3cos^2 x-sin^2 x = 0 3cos^2 x - (1-cos^2 x ) = 0 3cos^2 x -1+cos^2 x =0 cos^2 x = 1/4 cosx = 1/4 or cosx =-1/4